sir , please explain me the following question... S=1+3/1! +5/2! + 7/3!...............   options are 1) e 2)e-1 3) 3e 4) 2e …...... it’​s a kind of urgency........ please reply as soon as possible

Grade:12

1 Answers

Pratham Joshi
6 years ago
 
\sum2n+1/n(factorial)\doteq 2/(n-1)(factorial) + 1/n(factorial)= 2 + 2(e-1) + e= 3e